# Evaluating how LLM annotations represent diverse views on contentious topics

### v1

@@ Abstract
Researchers have proposed the use of generative large language models (LLMs) to label data for both research and applied settings. This literature emphasizes the improved performance of LLMs relative to other natural language models, noting that LLMs typically outperform other models on standard metrics such as accuracy, precision, recall, and F1 score. However, previous literature has also highlighted the bias embedded in language models, particularly around contentious topics such as potentially toxic content. This bias could result in labels applied by LLMs that disproportionately align with majority groups over a more diverse set of viewpoints. In this paper, we evaluate how LLMs represent diverse viewpoints on these contentious tasks. Across four annotation tasks on four datasets, we show that LLMs do not show substantial disagreement with annotators on the basis of demographics. Instead, the model, prompt, and disagreement between human annotators on the labeling task are far more predictive of LLM agreement. Our findings suggest that when using LLMs to annotate data, under-representing the views of particular groups is not a substantial concern. We conclude with a discussion of the implications for researchers and practitioners.
1 Introduction
@@
Recent research on large language models (LLMs) emphasizes their improved performance on a variety of natural language processing tasks including classification and text generation (Tan et al., 2024). In social science, researchers have proposed methods for using LLMs to label texts for concepts of interest in the social sciences (Rathje et al., 2024; Bail, 2024; Törnberg, 2023; Atreja et al., 2024). However, research on bias in machine learning systems shows that models frequently overrepresent the views of majority groups (Davani et al., 2022; Prabhakaran et al., 2021). The core concern motivating our paper is that by using these models for text annotation in social science research, such research will over-represent the views of majority groups and under-represent the views of minority groups.
@@
To evaluate this concern, we evaluate the performance of LLMs on four different text annotation tasks. We use the NLPositionality dataset (Santy et al., 2023), which contains labels on social media comments for toxicity; the POPQUORN offensiveness and politeness datasets (Pei & Jurgens, 2023), which contains labels on social media comments for offensiveness and labels on emails for politeness; and the Wikipedia comments dataset (Wulczyn et al., 2017), which contains labels on comments from Wikipedia edits for toxicity. We picked these datasets for two reasons. First, the underlying tasks are contentious in nature resulting in disagreement among human raters. Second, these datasets includes the demographics of annotators who generated the labels, allowing us to (a) evaluate the performance of LLMs on these tasks and (b) understand the extent to which LLMs agree with different demographic groups. We focus on gender, race, and education demographic groups because they were available in the datasets; prior work also acknowledges the potential bias related to these groups (e.g. see Thakur 2023; Omar et al. 2025; Kotek et al. 2023; 2024).
@@
For the LLMs, we use GPT-4o mini, Llama 3.3 70B Instruct, and Gemini 1.5 Flash. To ensure that our results are not prompt-specific, we also tested three different prompts for eliciting labels for each dataset. For each text from each dataset, we label the text using each of the three models and each of the three prompts, resulting in nine LLM labels per text. See Table 1 for a summary of these factors and their levels. We determine agreement by evaluating whether the model provides the same labels that the annotator did (e.g. if the annotator labeled the text toxic, and the model labeled the text toxic, then the model agrees with that annotator). We compute LLM’s agreement with each individual annotator and use logistic regression to understand the demographic factors that predict LLM agreement with human labelers.
Table 1: Factors and Levels in our Experiments
@@
Model GPT-4o mini Llama 3.3 70B Instruct Gemini 1.5 Flash
Dataset (task) NLPositionality (toxicity) POPQUORN Politeness (politeness) POPQUORN Offensiveness (offensiveness) Wikipedia comments (toxicity)
@@
Prompt 1 2 3
@@
Across all datasets, we find that LLMs do not show substantial disagreement with annotators on the basis of demographics. Rather, the difficulty of the labeling task (measured using disagreement among human annotators), the model used, and the prompt used are far more predictive of LLM agreement than the demographics of the annotators. Additionally, when LLMs tend to agree with one demographic group over the other, the demographic group is not consistent across datasets. For example, the LLMs tend to agree with white annotators more often on the offensiveness and politeness tasks, but tend to agree with non-white annotators more for the NLPositionality toxicity task. Within each task, there is no variation among models on the demographic groups with which the models tend to agree; GPT 4o-mini, Llama, and Gemini agree with the same demographic groups. These findings indicate that model agreement with demographic groups is dataset-specific and may be a function of the underlying population of annotators in the data generating process or of the annotation task itself rather than a function of underlying bias in the LLMs themselves.
2 Related Work
2.1 Bias in LLMs
@@
Previous research has focused on bias in natural language processing (NLP), highlighting how language models can manifest implicit biases existing in society (Bolukbasi et al., 2016; Caliskan et al., 2017; Garg et al., 2018; Zhao et al., 2019). Further, burgeoning research on LLMs shows that similar to their predecessors, LLMs exhibit biases related to gender, race, and other demographic factors (Kotek et al., 2024). Multiple studies have found gender bias to be prevalent in LLM outputs, manifesting in gendered word associations and biased narratives (Thakur, 2023; Omar et al., 2025) as well as gendered stereotypes (Kotek et al., 2023). Additional studies have found racial biases as well (Omar et al., 2025), demonstrating racist stereotypes or biased outcomes in medical LLMs (Hastings, 2024). LLMs exhibit intersectional biases in addition to racial and gender biases, with LLM-generated texts implying lower levels of agency for multiply marginalized demographic groups (e.g. Black females) (Wan & Chang, 2024). LLMs exhibit biases beyond just race and gender, including other protected groups such as sexuality and religion (Kotek et al., 2024). Moreover, LLMs may not only produce bias in the text they produce, but they may express even more bias when instructed to answer using personas (Dong et al., 2024).
2.2 Persona Effects in LLMs
@@
Persona effects are the ability of an LLM to embody a particular persona. For example, given a set of demographics, political beliefs, psychological profiles, or other information, LLMs can create responses aligning with individuals of those demographics. LLMs demonstrate consistent ability to manifest assigned personas, but their impact varies widely across different tasks and contexts, with some configurations of identities significantly affecting performance and others showing minimal influence. Hu & Collier (2024) found that personas account for less than 10% of the variance in data annotations, but including these annotations in LLM prompts significantly improves annotation accuracy. Persona assignment can shape LLM behavior in measurable ways. However, these benefits are not universal; prompting job-related identities to boost LLM accuracy did not show measurable improvements (Zheng et al., 2024). Persona assignment can also produce overly specific or stereotyped results. Gupta et al. (2023) showed that persona assignment used stereotypes when responding to questions after being assigned a persona, and Bisbee et al. (2024) shows that persona assignment to produce political opinions yields results with significantly less variation than the underlying population the LLM is attempting to represent. Taken together, the existing research on persona effects in LLMs shows that LLMs can accurately represent the viewpoints given a set of personas, but these viewpoints can also be biased with respect to the personas the researchers are aiming to represent.
2.3 LLMs for Text Labeling
@@
There is a growing body of research evaluating the efficacy of LLMs for text labeling tasks (Tan et al., 2024; Atreja et al., 2024). Many of these studies find that LLMs exhibit higher performance in classification tasks, and LLMs can be cheaper than hiring human annotators (Li et al., 2023; Törnberg, 2023). Moreover, there are particular tasks where LLM labeling can yield pro-social results—for example, using LLMs to label toxic online speech rather than exposing moderators to harmful online content (Li et al., 2024). Other research evaluating the use of LLMs in social research shows that LLMs display significant persona effects, meaning when prompted, they can accurately represent the opinions of the persona prompted (Hu & Collier, 2024). However, this accuracy comes at a cost; persona prompting can be overly precise, masking variance among the demographic groups they aim to represent (Bisbee et al., 2024; Hu & Collier, 2024). Among subjective annotation tasks, which are often the subject of papers proposing the use of large language models (e.g. Törnberg 2023), it is unclear whether LLMs are biased in their annotations.
2.4 Annotators and Subjective Annotation Tasks
@@
Models are frequently trained on data labeled by annotators. For subjective annotation tasks–i.e. tasks where the label may not have a ground truth and instead rely on the subjective judgments of the data annotators–data annotations can replicate the same biases that exist in humans (Hube et al., 2019; Díaz, 2020). Recent research on subjective data annotations shows that annotator characteristics influence task outcomes (Al Kuwatly et al., 2020; Hube et al., 2019; Díaz, 2020). Aggregating over annotators to determine a ground truth label (for example, by taking the majority label across all annotators) can systematically under-represent the views of minority annotators (Davani et al., 2022; Prabhakaran et al., 2021). The pool of annotators—their demographics, cultural backgrounds, political beliefs, and personality traits—and the aggregation decisions made by researchers can create biased training data. Machine learning models are then trained on these biased annotations, manifesting bias in model outputs. Díaz (2020) shows how annotator bias in subjective annotation tasks biases downstream model predictions.
@@
Disagreement in annotations is not necessarily a result of poor performance on the task, but can be the consequence of different subjective beliefs (Wan et al., 2023). Using LLMs to label texts essentially treats an LLM as a data annotator, which may suffer from the same biases exhibited when aggregating over multiple annotators since LLMs can be understood as aggregations over large amounts of training data.
3 Data & Methods
@@
To account for the diversity of language tasks researchers may use LLMs for, we use four open datasets covering four annotation tasks in this study (Santy et al., 2023; Pei & Jurgens, 2023; Wulczyn et al., 2017). These datasets are labeled for contentious topics, which are the subject of much of the research related to LLM annotations due to the potential for reducing human annotators’ exposure to harmful content (Li et al., 2024). These datasets are also among the few datasets in NLP that include the demographics of annotators, allowing us to understand how LLM annotations represent (or do not represent) the views of different annotators. We describe the datasets below and outline our analytical method for evaluating LLM annotation biases. In Table 2 we show a summary of the demographics in each dataset. For education, the majority group is non-college educated, and the minority group is college educated; for gender, the majority group is men, and the minority group is non-men; and for race, the majority group is white and the minority group is non-white.
3.1 NLPositionality Toxicity Data
@@
The NLPositionality Toxicity Data is a collection of social media posts that contain binary labels indicating whether or not those posts are toxic (Santy et al., 2023). Annotators were sourced through Lab in the Wild, an opt-in crowd-sourcing platform to evaluate the alignment of models with different demographic groups. From these data, we used demographic data for gender, race, and education. For the annotators, there are 336 annotators that identify as men and 711 annotators that identify as non-men; 296 white annotators and 766 non-white annotators; and 253 college-educated annotators and 614 non-college educated annotators. There are 299 unique texts annotated by 1,082 unique annotators.
3.2 POPQUORN Offensiveness Data
@@
The POPQUORN Offensiveness dataset contains social media comments with ordinal labels on a scale of 1 to 5 (1 being not offensive at all and 5 being very offensive). The annotators are sourced through a quota-sampled pool of workers on Lucid with the aim of being generally representative of the U.S. population. From these data, we use gender, race, and education demographics. There are 262 unique annotators of which 124 identify as men (138 non-men), 197 identify as white (65 non-white), and 171 college educated (85 non-college educated). The POPQUORN Offensiveness dataset consists of 1,500 unique comments randomly sampled from the Ruddit dataset (Hada et al., 2021).
3.3 POPQUORN Politeness Data
@@
The POPQUORN Politeness data set contains emails with ordinal labels on a scale of 1 to 5 (1 being not polite at all and 5 being very polite). The annotators are sourced through a quota-sampled pool of workers on Lucid with the aim of being generally representative of the U.S. population, though from a different pool than the POPQUORN Offensiveness dataset. From these data, we use gender, race, and education demographics. There are 506 unique annotators with 236 identifying as men, 267 identifying as non-men, 367 identifying as white, 135 identifying as non-white, 334 college educated, and 155 non-college educated. The dataset contains 3,718 unique emails from Shetty & Adibi (2004).
3.4 Wikipedia Toxicity Data
@@
The Wikipedia toxicity data set contains comments in Wikipedia edits which contains binary labels for comments indicating whether or not they are toxic (Wulczyn et al., 2017). The annotators were sourced on Crowdflower. From these data, we use gender and education demographics; this dataset does not contain information about annotators’ race. The dataset consists of 2,893 unique annotators, of which 1,913 identify as men, 980 identify as non-men, 1,980 are college educated, and 911 are not college educated. The dataset consists of 159,463 unique texts of which we sample 2,095 to label. The details of this sampling strategy can be found below.
Table 2: Demographics of annotators across datasets
@@
POPQUORN Offensivness Education 171 85 6
POPQUORN Offensivness Gender 138 124 0
POPQUORN Offensivness Race/Ethnicity 65 197 0
POPQUORN Politeness Education 334 155 17
POPQUORN Politeness Gender 267 236 3
POPQUORN Politeness Race/Ethnicity 135 367 4
@@
3.5 Data Processing and Sampling

### v2

@@ Abstract
Researchers have proposed the use of generative large language models (LLMs) to label data for research and applied settings. This literature emphasizes the improved performance of these models relative to other natural language models, noting that generative LLMs typically outperform other models and even humans across several metrics. Previous literature has examined bias across many applications and contexts, but less work has focused specifically on bias in generative LLMs’ responses to subjective annotation tasks. This bias could result in labels applied by LLMs that disproportionately align with majority groups over a more diverse set of viewpoints. In this paper, we evaluate how LLMs represent diverse viewpoints on these contentious tasks. Across four annotation tasks on four datasets, we show that LLMs do not show systematic substantial disagreement with annotators on the basis of demographics. Rather, we find that multiple LLMs tend to be biased in the same directions on the same demographic categories within the same datasets. Moreover, the disagreement between human annotators on the labeling task—a measure of item difficulty—is far more predictive of LLM agreement with human annotators. We conclude with a discussion of the implications for researchers and practitioners using LLMs for automated data annotation tasks. Specifically, we emphasize that fairness evaluations must be contextual, model choice alone will not solve potential issues of bias, and item difficulty must be integrated into bias assessments.
Index Terms:
large language models, data annotation, text mining, bias, fairness
I Introduction
@@
Recent research on large language models (LLMs) emphasizes their improved performance in a variety of natural language processing tasks, including classification and text generation [1]. In the social sciences, researchers have proposed using generative LLMs to label texts for concepts of interest, usually in zero-shot settings [2, 3, 4, 5, 6]. While research shows that machine learning models frequently overrepresent the views of majority groups (see, e.g., [7, 8]), this has not been systematically studied for generative LLMs. The core concern motivating our paper is that by using these models for text annotation in social science research, such research will over-represent the views of majority groups and under-represent the views of minority groups.
@@
To better understand this concern, we propose systematically evaluating the performance of a suite of generative LLMs on four different text annotation tasks. We use the NLPositionality dataset [9], which contains labels on social media comments for toxicity; the POPQUORN offensiveness and politeness datasets [10], which contains labels on social media comments for offensiveness and labels on emails for politeness; and the Wikipedia comments dataset [11], which contains labels on comments from Wikipedia edits for toxicity. We selected these datasets because (1) the underlying tasks are subjective and often contentious in nature, resulting in disagreement among human raters; and (2) these datasets includes the demographics of annotators who provided the labels, allowing us to evaluate the performance of LLMs on these tasks and understand the extent to which LLMs agree with different demographic groups. We focus specifically on gender, race, and education because they were available in all the datasets; prior work also acknowledges the potential bias related to these groups (see, e.g., [12, 13, 14, 15]).
@@
For LLMs, we use GPT-4o mini, Llama 3.3 70B Instruct, and Gemini 1.5 Flash. These are three of the most commonly used LLMs in social science research. To ensure that our results are not prompt-specific, we also tested three semantically equivalent prompts for generating labels for each dataset. We label each text from each dataset using each of the three models and each of the three semantically equivalent prompts, resulting in nine LLM labels per text. See Table I for a summary of these factors and their levels. We determine agreement by evaluating whether the model generates the same label that the annotator did (e.g., if the annotator labeled the text toxic and the model labeled the text toxic, then the model agrees with that annotator). We compute the LLMs’ agreements with each individual annotator and use logistic regression to understand the demographic factors that predict LLM agreement with human annotators.
TABLE I: Factors and Levels in our Experiments
@@
Models GPT-4o mini Llama 3.3 70B Instruct Gemini 1.5 Flash
Datasets (tasks) NLPositionality (toxicity) POPQUORN Politeness (politeness) POPQUORN Offensiveness (offensiveness) Wikipedia comments (toxicity)
@@
Prompts Prompt Version 1 Prompt Version 2 Prompt Version 3
@@
Across all datasets, we find that LLMs do not show systematic and substantial disagreement with annotators on the basis of demographics. While there are statistically significant biases, the direction of the bias of each demographic category is generally the same across all LLMs. Additionally, when LLMs tend to agree with one demographic group over the other, the demographic group is not consistent across datasets. In other words, biases tend to be dataset-specific rather than LLM-specific. For example, the LLMs tend to agree with white annotators more often on the offensiveness and politeness tasks, but tend to agree more with non-white annotators on the NLPositionality toxicity task. Within each task, there is no variation among the models on the demographic groups with which the models tend to agree: GPT 4o-mini, Llama, and Gemini agree with the same demographic groups. Moreover, the difficulty of the labeling task (measured using disagreement among human annotators) is much more predictive of LLM agreement than the demographics of the annotators. Our findings indicate that LLM agreement with specific demographic groups may be a function of the data and how it interacts with the LLM’s pretrained embedded information, rather than a function of systematic biases of the LLM across all subjective annotation tasks.
In summary, our contributions are as follows.
  • We provide the first systematic evaluation of demographic bias in LLMs when using LLMs for social science annotation tasks. We use four datasets, three popular LLMs, and three semantically equivalent prompts for each dataset in our analysis. Beyond analyzing demographic features, we also explicitly model item difficulty using entropy.
  • We find that LLMs do not show systematic disagreement with annotators on demographics (specifically, gender, race, and education). We find that all LLMs tend to be biased in the same directions on the same dataset. Moreover, the difficulty of the labeling task is most predictive of LLM agreement with human annotators.
  • We discuss the implications for researchers and practitioners interested in using generative LLMs for automated data annotation tasks. Most importantly, we emphasize that fairness evaluations should be context-specific, model choice alone will not solve bias, and item difficulty must be integrated into bias assessments.
II Related Work
II-A Bias in LLMs
@@
Previous research has focused on bias in natural language processing, highlighting how embeddings and language models can embed and manifest implicit biases existing in society [16, 17, 18, 19]. Further, burgeoning research on LLMs shows that similar to their predecessors, LLMs exhibit biases related to gender, race, and other demographic factors on a range of tasks such as job recommendations, [20], healthcare [21], and occupation prediction [15]. Multiple studies have found that gender bias is prevalent in LLM outputs, manifesting in gendered word associations and biased narratives [12, 13] as well as gendered stereotypes [14]. Additional studies have also found racial biases [13], such as medical LLMs generating racist stereotypes or biased outcomes [22]. LLMs exhibit intersectional biases in addition to racial and gender biases, with LLM-generated texts implying lower levels of agency for multiply marginalized demographic groups (e.g., Black females) [23]. LLMs also exhibit biases beyond just race and gender, including other protected groups such as sexuality and religion [15]. Studies have also found that LLMs exhibit sociodemographic biases even when they decline to respond [24]. Moreover, LLMs may not only produce bias in the text they produce, but may express even greater bias when instructed to answer using personas [25]. Our work builds on this literature by specifically looking at bias in subjective social science annotation tasks using generative LLMs.
II-B Persona Effects in LLMs
@@
Persona effects are the ability of an LLM to embody a particular persona. For example, given a set of demographics, political beliefs, psychological profiles, or other information, LLMs can create responses that align with individuals of those demographics. LLMs demonstrate consistent ability to manifest assigned personas, but their impact varies widely across different tasks and contexts, with some configurations of identities significantly affecting performance and others showing minimal influence. The authors of [26] found that personas account for less than 10% of the variance in data annotations, but including these annotations in LLM prompts significantly improves annotation accuracy; in other words, persona assignment can shape LLM behaviors in measurable ways. However, these benefits are not universal. Prompting job-related identities did not show measurable improvements [27]. Persona assignment can also produce overly specific or stereotyped results. The authors of [28] showed that persona assignment used stereotypes when responding to questions after being assigned a persona, and the authors of [29] showed that persona assignment to produce political opinions yields results with significantly less variation than the underlying population the LLM is attempting to represent. Taken together, the existing research on persona effects in LLMs shows that LLMs can accurately represent the viewpoints given a set of personas, but these viewpoints can also be biased with respect to the personas the researchers are aiming to represent. Our work does not use prompts that induce a persona. We use three semantically equivalent prompts that do not contain any demographic information. Our aim is to better understand if the LLM’s generated response to identity-free prompts more strongly correlates with specific demographic categories.
II-C LLMs for Text Labeling
@@
There is a growing body of research evaluating the efficacy of generative LLMs for text labeling tasks (see, e.g., [1, 5]). Many of these studies find that LLMs exhibit higher performance in classification tasks, and LLMs can be cheaper than hiring human annotators [30, 31, 32]. Moreover, there are particular tasks where LLM labeling can yield pro-social results, such as using LLMs to label toxic online speech rather than exposing moderators to harmful online content [33]. Among subjective annotation tasks, which are often the subject of papers proposing the use of LLMs (e.g., [32]), it is unclear whether LLMs are biased in their annotations. Our paper aims to analyze whether LLMs are biased in their generated labels on subjective annotation tasks.
II-D Annotators and Subjective Annotation Tasks
@@
Models are frequently trained on data labeled by annotators. For subjective annotation tasks—tasks where the label may not have a ground truth and instead rely on the subjective judgments of the data annotators—data annotations can replicate the same biases that exist in humans [34, 35, 36]. Recent research on subjective data annotations shows that annotator characteristics influence task outcomes [37, 34, 35]. Aggregating over annotators to determine a ground truth label (for example, by taking the majority label across all annotators) can systematically underrepresent the views of minority annotators [7, 8]. The pool of annotators—their demographics, cultural backgrounds, political beliefs, and personality traits—and the aggregation decisions made by researchers can create biased training data. Machine learning models are then trained on these biased annotations, manifesting bias in model outputs. [35] shows how annotator bias in subjective annotation tasks biases downstream model predictions.
@@
Disagreement in annotations is not necessarily a result of poor performance on the task, but can be the consequence of different subjective beliefs [38]. Using LLMs to label texts essentially treats an LLM as a data annotator, which may suffer from the same biases exhibited when aggregating over multiple annotators since LLMs can be understood as aggregations over large amounts of training data. We aim to analyze this type of bias that LLMs may exhibit on subjective annotation tasks.
To combat this, some works have attempted to build into their classification pipelines a diverse set of perspectives, often called perspective-aware models (see, e.g., [39, 40, 41]). However, our work focuses exclusively on using generative LLMs without any additional gradient-based updates or architectural modifications to the models. Our prompt-based approach largely aligns with how social scientists typically use these models.
The work closest to ours is [42]. They also analyzed the POPQUORN dataset’s offensiveness and politeness datasets to understand how LLM annotations may differentially correlate with specific demographic categories. Their findings on this dataset are similar to ours. In our work, however, we looked at two additional datasets, allowing us to better understand if the same patterns generalize outside the POPQUORN suite of datasets. We also explicitly modeled text classification difficulty using an entropy approach, while they used random effects for each item.
III Data & Methods
@@
To account for the diversity of language tasks researchers may use LLMs for, we used four open datasets covering four annotation tasks in this study [9, 10, 11]. These datasets are labeled for contentious topics, which are the subject of much research related to LLM annotations due to the potential for reducing human annotators’ exposure to harmful content [33]. These datasets are also among the few in NLP that include the demographics of annotators, allowing us to understand how LLM annotations represent (or do not represent) the views of different annotators. We describe the datasets below and outline our analytical method for evaluating LLM annotation biases. Table II shows a summary of the demographics in each dataset. For education, the majority group is non-college educated and the minority group is college educated; for gender, the majority group is men and the minority group is non-men; and for race, the majority group is white and the minority group is non-white.
III-A NLPositionality Toxicity Data
@@
The NLPositionality Toxicity Data is a collection of social media posts that contain binary labels indicating whether or not those posts are toxic [9]. Annotators were sourced through Lab in the Wild, an opt-in crowd-sourcing platform to evaluate the alignment of models with different demographic groups. From this dataset, we used gender, race, and education demographics. There are 299 unique texts annotated by 1,082 unique annotators.
III-B POPQUORN Offensiveness Data
@@
The POPQUORN Offensiveness dataset contains social media comments with ordinal labels on a scale of 1 to 5 (1 being not offensive at all and 5 being very offensive). The annotators are sourced through a quota-sampled pool of workers on Lucid with the aim of being generally representative of the U.S. population. From this dataset, we used gender, race, and education demographics. The POPQUORN Offensiveness dataset consists of 1,500 unique comments randomly sampled from the Ruddit dataset [43].
III-C POPQUORN Politeness Data
@@
The POPQUORN Politeness data set contains emails with ordinal labels on a scale of 1 to 5 (1 being not polite at all and 5 being very polite). The annotators are sourced through a quota-sampled pool of workers on Lucid with the aim of being generally representative of the U.S. population, though from a different pool than the POPQUORN Offensiveness dataset. From this dataset, we used gender, race, and education demographics. The dataset contains 3,718 unique emails from [44].
III-D Wikipedia Toxicity Data
@@
The Wikipedia toxicity data set contains comments in Wikipedia edits which contains binary labels for comments indicating whether or not they are toxic [11]. The annotators were sourced on CrowdFlower. From this dataset, we used gender and education demographics; this dataset does not contain information about annotators’ race. The dataset consists of 159,463 unique texts. Due to budget constraints, we used stratified samples of this dataset. We stratified this sample based on label entropy, ensuring that our sample is balanced across texts of varying levels of disagreement among annotators. This sampling process yielded 2,095 unique texts.
TABLE II: Demographics of annotators across datasets. POPQUORN is abbreviated as PQ.
@@
PQ Offensivness Education 171 85 6
PQ Offensivness Gender 138 124 0
PQ Offensivness Race/Ethnicity 65 197 0
PQ Politeness Education 334 155 17
PQ Politeness Gender 267 236 3
PQ Politeness Race/Ethnicity 135 367 4
@@
III-E Data Processing and Sampling